OpenStudy (anonymous):

Please help, I will medal and fan 4. 3+d<3-d 5. The sum of two consecutive integers is at least 46. What is the least possible pair of integers? 21 and 23 23 and 24 22 and 23 24 and 25

OpenStudy (abdullahm):

3+d < 3-d Add d on both sides Tell me what you get :)

OpenStudy (abdullahm):

Basically think of the inequality as an equal sign :)

OpenStudy (anonymous):

d>6?

OpenStudy (abdullahm):

Well let me see... mhmm... 3 + d < 3 - d 3 + 2d < 3 Subtract 3 on both sides 2d < 0 Divide by 2 on both sides d < 0

OpenStudy (abdullahm):

So that's wrong :P

OpenStudy (anonymous):

oh, ok I see what you did.

OpenStudy (anonymous):

can you help me with number 5?

OpenStudy (abdullahm):

sure :)

OpenStudy (abdullahm):

Consecutive integers means like back to back, so like x and (x + 1)

OpenStudy (abdullahm):

and the sum of the two consecutive integers is at least 46 so... x + x + 1 \(\ge\) 46 2x + 1 \(\ge\) 46 Solve for x :)

OpenStudy (anonymous):

2x + 47 dividing 2x on both sides would give me 23.5

OpenStudy (abdullahm):

So that means that both numbers over 2 = 23.5

OpenStudy (abdullahm):

That would mean if one number is 23 and the other is 24, (23+24)/2 = 23.5

OpenStudy (anonymous):

so b

OpenStudy (abdullahm):

yes

OpenStudy (anonymous):

Thanks so much! here is you medal and fan

OpenStudy (abdullahm):

No problem (: